# AcE: Attribution-Controlled Knowledge Editing for Multi-hop Factual Recall

### v1

@@ 2 Related Work
Effective knowledge editing depends on understanding how LLMs store information. FFN layers have been shown to act as key-value stores, with values encoding semantic content (Geva et al., 2020), and recent work has localized factual knowledge at finer granularity (Dai et al., 2021; Hernandez et al., 2023). Yu & Ananiadou (2023) revealed that predictions are driven by interactions between query and value neurons, with value neurons exhibiting consistent relational semantics. Building on this, AcE employs a neuron-level approach to trace and modulate these activation pathways, enabling targeted and interpretable editing in multi-hop reasoning.
@@ 3.1 Modeling and Task
Following the same notation as in (Yu & Ananiadou, 2023), $fc2_{k}^{l}$ is the subvalue of FFN which is the column kthkth of $W_{fc2}^{l}$, and its coefficient score $m_{i,k}^{l}$ is calculated by residual output $h^{l-1}_{i}+A^{l}_{i}$ and FFN subkey $fc1_{k}^{l}$, the kthk-th row of $W_{fc1}^{l}$. Similarly, the attention output $A_{i}^{l}$ can be represented as:
@@ 3.2 Attribution Metrics
To jointly account for both the variable neuron $v$ and the conditioning variable $x$ in the probabilistic framework. Based on Yu & Ananiadou (2023), we find log probability increase could efficiently evaluate the model’s distribution change by a neuron. Define the log probability increase as importance score $\mathcal{I}$ for vectors, which satisfies $\mathcal{I}(x+v)\approx\mathcal{I}(x)+\mathcal{I}(v)$. If $v^{l}$ is a vector in lthlth attention layer, the importance score of $v^{l}$ is:
@@ 4 Mechanism of Multi-hop Reasoning
How do language models store and retrieve knowledge when performing complex multi-hop reasoning? We begin by examining how semantically related knowledge is organized within transformer components (Section 4.1), revealing consistent patterns that challenge conventional wisdom about knowledge storage. Building on these structural insights, we then trace how information propagates through reasoning chains (Section 4.2), uncovering a sophisticated query-value activation mechanism that progressively accumulates evidence toward final answers. Our analysis spans both GPT-J (Wang & Komatsuzaki, 2021) and the more advanced Qwen3-8B (Team, 2025), with the latter exhibiting even more fine-grained activation patterns that offer new insights into reasoning dynamics.
@@ 4.2 How is the Information Accumulated?
To validate the functional importance of these patterns, we conducted targeted interventions on query neurons. When we ablated the top 100 query neurons from two layers showing peak activation for 2-hop requests ($f_{q16}$ and $f_{q18}$), model capability decreased by 46.2% and 61.9%, respectively. Moreover, the number of activated value neurons in subsequent layers ($f_{17,18,19}$) dropped dramatically from (28,16,33) to (6,4,7), while other value neurons showed minimal change (12 neurons decreased). We show more details on Qwen3 while we explore the forward processes in Appendix E.
@@ 5 AcE: Attribution-Controlled Knowledge Editing
Building upon identified critical layers, we conducted sequential knowledge editing to demonstrate the model’s enhanced knowledge acquisition through intensified activation patterns. Based on previous locate-then-edit paradigm (Li et al., 2024), we apply editing on FFNs and keep attention heads unchanged, while the general semantic information saved in attention heads should not be changed. For $l$th layer FFN, it’s output of the $i$th token ’s would be $W_{fc2}^{l}\sigma(W_{fc2}^{l}h^{l-1}_{i})$, and there’s no attention input in GPT-J, where $\sigma$ is the non-linear activation function. $\sigma(W_{fc2}^{l}h^{l-1}_{i})$ take the responsibility as keys, denoted as $k_{i}$. Whole FFN output is the value $v_{i}=W_{fc2}^{l}\sigma(W_{fc2}^{l}h^{l-1}_{i})$, whose subvalue matrix $W_{fc2}^{l}$ denotes the weight of the model’s values needs to be modified. So we aim to modify subvalue matrix $W_{fc2}^{l}$ s.t. $W_{fc2}^{l}k=v^{*}$, where $v^{*}$ represents the new values (factual knowledge). We use backbone PMET (Li et al., 2024) to complete editing process, details shown in Algorithm C.
@@ 6.1 Experimental Setup
MQuAKE-3K (Zhong et al., 2023) is a benchmark dataset for evaluating large language models’ multi-hop fact recall capability after knowledge editing. Each instance contains a multi-hop fact recall chain (interconnected triples) and corresponding textual questions, requiring models to maintain coherent reasoning after applying single-hop knowledge edits via edit prompts. This design simulates cascading effects of real-world knowledge updates, providing a systematic framework to assess models’ dynamic knowledge management.
To faithfully evaluate the model’s capabilities and assess its capacity to assimilate edited knowledge, we applied filtering to the original dataset, resulting in a curated subset where the base model consistently achieves accurate reasoning.
@@
We apply our AcE on the model GPT-J (6B) and Qwen3-8B. We use Multi-hop requests answering accuracy as the metric to evaluate the performance of our edited model. We use PMET as our model’s primary backbone for editing. More settings details see in Appendix F.
@@ 6.2 Main Results
Table 2 demonstrates the general performance of various established methods and AcE under the different classes of data settings on our subset of MQuAKE-3K. To enable precise identification of critical layers store the knowledge and activation patterns during model editing, we systematically integrated in-context priming and Chain-of-Thought reasoning across all editing prompts. This dual-strategy architecture ensures optimal knowledge editing efficacy through targeted activation within the model’s parameter space.
# Edits refers to the number of how many individual facts in the reasoning chain are edits, its maximum equals to the number of hops in the dataset. As evidenced in Table 2, the proposed AcE framework demonstrates consistent superiority over existing methodologies across various evaluation metrics. Our method outperformances $9.44\%$ and $37.46\%$ on GPT-J and Qwen3-8B. Traditional Locate-then Edit paradigm performances even worse on Qwen3-8B due to the fixed editing positions, and AcE shows much flexibility on reasoning models due to the unfixed q-v positions as discussed in Section 6.4.
@@ 6.4 Analysis
Now we can explain why the existing KE methods failed in multi-hop reasoning. Existing KE methods overlooked the value layers in deeper locations, even regarded the output generation layers as the value layers, where the knowledge truly be stored in latter ones (details in Appendix E). And the worst is, these KE methods ignored the importance of editing the query layers.
@@
Case Study. As a case study, we examined the residual stream for the single-hop request: “Tim Duncan plays the sport of”. The model exhibited the largest increase in importance at tokens where semantics converge, such as “plays” and “of”, with reaching $0.9932$ and $0.8469$, respectively. The top predicted tokens were highly interpretable (e.g., “basketball”, “ball”, “NBA”). Meanwhile, at other transitional tokens, the model demonstrated stronger exploratory capabilities by predicting incoherent tokens. We select 27 neurons whose associated vocabulary included the correct target token and performed 1000 sampling trials with these neurons removed. The model’s accuracy under this condition dropped to merely 3.2%, indicating the crucial role these interpretable neurons play in generating correct responses. In contrast, when we ablated 27 neurons with high importance scores but lacking correct semantic interpretability, the model maintained 59.4% accuracy. This suggests the final generation of correct output tokens depends on a sparse set of highly specialized, interpretable neurons.We also observe some neurons serve as q-v shared neurons at the same time, which are highly interpretable, the details of this residual stream see in Appendix D.
@@ 7 Conclusion
Appendix A Related Work in Details
@@
The work by Yu & Ananiadou (2023) provides a crucial foundation for our methodology. They demonstrated that both attention and FFN layers operate via query neurons that activate specific value neurons to produce final predictions. These value neurons were found to be semantically interpretable when projected into the vocabulary space, and similar types of relations (e.g., “capital of X,” “birthplace of Y”) were shown to activate neurons in structurally consistent locations across different subjects. While IFMET focused on attribution and interpretation, AcE leverages these neuron-level insights to develop a more targeted and mechanistically grounded KE strategy, particularly for the complex dynamics of multi-hop reasoning where intermediate implicit subjects act as activators.
@@
Appendix B Subset of MQuAKE Dataset
B.1 Subset Construction
@@
B.2 Dataset Labeling
We use GPT-4o to label the knowledge in the dataset. We show the prompts we used in Appendix G.3. We labeled the dataset related to the knowledge, which are Geographic location, Organization, Personal attributes, Sports, Entertainment, Language and culture, Education, Religion, Literature and Event, which consists of more detailed smaller classes.
Appendix C AcE Optimization
@@
Appendix D Details of Internal Fine-Grained Reasoning Q-V Pairs Flow in Qwen3-8B
@@
Appendix E More Details of the Analysis
@@
Appendix F Experimental Settings
@@
In our edits, our configuration for PMET adheres to the settings specified by (Li et al., 2024). Initially, we set $\varphi=1$ and $0\leq\mu\leq 1$ to manage the retention of the model’s original knowledge. As $\mu$ increases, the retention level also increases, while $\varphi$ exhibits the opposite trend. After maximizing the probability of the target knowledge, we reduce $\varphi$ to 0.1 to preserve the original knowledge as much as possible. Optimization is halted when $D_{KL}<0.01$. On GPT-J and Qwen3-8B, for estimating the covariance matrix (i.e., the set of previously memorized keys $C_{0}$), we sample 10,0000 times on Wikitext in fp32 precision and set $\lambda=6000$. When optimizing, we limit the total optimization steps to 30 with a learning rate of 0.2. All our experiments were conducted using the MQuAKE dataset. To test the accuracy of answers to multi-hop questions, we adhered to the few-shot and Chain of Thought (CoT) templates in Appendix G.1 and procedures.
Appendix G Prompts
@@
G.1 Few-shot and Chain-of-Thought Evaluation Prompts

### v2

@@ 2 Related Work
Effective knowledge editing depends on understanding how LLMs store information. FFN layers have been shown to act as key-value stores, with values encoding semantic content (Geva et al., 2020), and recent work has localized factual knowledge at finer granularity (Dai et al., 2021; Hernandez et al., 2023). Yu and Ananiadou (2023) revealed that predictions are driven by interactions between query and value neurons, with value neurons exhibiting consistent relational semantics. Building on this, AcE employs a neuron-level approach to trace and modulate these activation pathways, enabling targeted and interpretable editing in multi-hop reasoning.
@@ 3.1 Modeling and Task
Following the same notation as in (Yu and Ananiadou, 2023), $fc2_{k}^{l}$ is the subvalue of FFN which is the column kthkth of $W_{fc2}^{l}$, and its coefficient score $m_{i,k}^{l}$ is calculated by residual output $h^{l-1}_{i}+A^{l}_{i}$ and FFN subkey $fc1_{k}^{l}$, the kthk-th row of $W_{fc1}^{l}$. Similarly, the attention output $A_{i}^{l}$ can be represented as:
@@ 3.2 Attribution Metrics
To jointly account for both the variable neuron $v$ and the conditioning variable $x$ in the probabilistic framework. Based on Yu and Ananiadou (2023), we find log probability increase could efficiently evaluate the model’s distribution change by a neuron. Define the log probability increase as importance score $\mathcal{I}$ for vectors, which satisfies $\mathcal{I}(x+v)\approx\mathcal{I}(x)+\mathcal{I}(v)$. If $v^{l}$ is a vector in lthlth attention layer, the importance score of $v^{l}$ is:
@@ 4 Mechanism of Multi-hop Reasoning
How do language models store and retrieve knowledge when performing complex multi-hop reasoning? We begin by examining how semantically related knowledge is organized within transformer components (Section 4.1), revealing consistent patterns that challenge conventional wisdom about knowledge storage. Building on these structural insights, we then trace how information propagates through reasoning chains (Section 4.2), uncovering a sophisticated query-value activation mechanism that progressively accumulates evidence toward final answers. Our analysis spans both GPT-J (Wang and Komatsuzaki, 2021) and the more advanced Qwen3-8B (Team, 2025), with the latter exhibiting even more fine-grained activation patterns that offer new insights into reasoning dynamics.
@@ 4.2 How is the Information Accumulated?
To validate the functional importance of these patterns, we conducted targeted interventions on query neurons. When we ablated the top 100 query neurons from two layers showing peak activation for 2-hop requests ($f_{q16}$ and $f_{q18}$), model capability decreased by 46.2% and 61.9%, respectively. Moreover, the number of activated value neurons in subsequent layers ($f_{17,18,19}$) dropped dramatically from (28,16,33) to (6,4,7), while other value neurons showed minimal change (12 neurons decreased). We show more details on Qwen3 while we explore the forward processes in Appendix F.
@@ 5 AcE: Attribution-Controlled Knowledge Editing
Building upon identified critical layers, we conducted sequential knowledge editing to demonstrate the model’s enhanced knowledge acquisition through intensified activation patterns. Based on previous locate-then-edit paradigm (Li et al., 2024), we apply editing on FFNs and keep attention heads unchanged, while the general semantic information saved in attention heads should not be changed. For $l$th layer FFN, it’s output of the $i$th token ’s would be $W_{fc2}^{l}\sigma(W_{fc2}^{l}h^{l-1}_{i})$, and there’s no attention input in GPT-J, where $\sigma$ is the non-linear activation function. $\sigma(W_{fc2}^{l}h^{l-1}_{i})$ take the responsibility as keys, denoted as $k_{i}$. Whole FFN output is the value $v_{i}=W_{fc2}^{l}\sigma(W_{fc2}^{l}h^{l-1}_{i})$, whose subvalue matrix $W_{fc2}^{l}$ denotes the weight of the model’s values needs to be modified. So we aim to modify subvalue matrix $W_{fc2}^{l}$ s.t. $W_{fc2}^{l}k=v^{*}$, where $v^{*}$ represents the new values (factual knowledge). We use backbone PMET (Li et al., 2024) to complete editing process, details shown in Algorithm D.
@@ 6.1 Experimental Setup
MQuAKE-3K (Zhong et al., 2023) is a benchmark dataset for evaluating large language models’ multi-hop fact recall capability after knowledge editing. Each instance contains a multi-hop fact recall chain (interconnected triples) and corresponding textual questions, requiring models to maintain coherent reasoning after applying single-hop knowledge edits via edit prompts. This design simulates cascading effects of real-world knowledge updates, providing a systematic framework to assess models’ dynamic knowledge management. To faithfully evaluate the model’s capabilities and assess its capacity to assimilate edited knowledge, we applied filtering to the original dataset, resulting in a curated subset where the base model consistently achieves accurate reasoning.
@@
We apply our AcE on the model GPT-J (6B) and Qwen3-8B. We use Multi-hop requests answering accuracy as the metric to evaluate the performance of our edited model. We use PMET as our model’s primary backbone for editing. More settings details see in Appendix G.
@@ 6.2 Main Results
Table 2 demonstrates the general performance of various established methods and AcE under the different classes of data settings on our subset of MQuAKE-3K. To enable precise identification of critical layers store the knowledge and activation patterns during model editing, we systematically integrated in-context priming and Chain-of-Thought reasoning across all editing prompts. This dual-strategy architecture ensures optimal knowledge editing efficacy through targeted activation within the model’s parameter space. # Edits refers to the number of how many individual facts in the reasoning chain are edits, its maximum equals to the number of hops in the dataset. As evidenced in Table 2, the proposed AcE framework demonstrates consistent superiority over existing methodologies across various evaluation metrics. Our method outperformances $9.44\%$ and $37.46\%$ on GPT-J and Qwen3-8B. Traditional Locate-then Edit paradigm performances even worse on Qwen3-8B due to the fixed editing positions, and AcE shows much flexibility on reasoning models due to the unfixed q-v positions as discussed in Section 6.4.
@@ 6.3 Ablation Study
Few-Shots Prompts.
Selecting appropriate Few-Shot prompts for in-context learning within the locate-then-edit paradigm enables the model to learn from challenging multi-hop reasoning editing examples, thereby improving its ability to answer questions accurately. We conducted ablative experiments on the prompts employed in our study. We conducted Zero-Shot, One-Shot and OOD Few-Shots (no overlap between evaluation set and prompts) prompts when evaluating.
As shown in Table 5, even with less information, the performance degradation is only 9.4% (Zero-Shot, GPT-J), 5.7% (One-Shot, GPT-J), 0.4% (OOD Few-Shots, GPT-J), 9.5% (Zero-Shot, Qwen3), 3.1% (One-Shot, Qwen3), and 0.6% (OOD Few-Shots, Qwen3) compared to the original. This demonstrates that AcE maintains strong robustness under reduced in-context learning information, providing evidence that the effectiveness of knowledge editing is inherent rather than reliant on in-context learning.
Table 5: The results of ablation experiments on different CoT Prompts upon GPT-J and Qwen3.
Editor Avg.
GPT-J (Zero-Shot) $42.08$
GPT-J (One-Shot) $43.79$
GPT-J (OOD Few-Shots) $46.27$
Qwen3 (Zero-Shot) $53.19$
Qwen3 (One-Shot) $56.46$
Qwen3 (OOD Few-Shots) $57.92$
@@ 6.4 Analysis
Now we can explain why the existing KE methods failed in multi-hop reasoning. Existing KE methods overlooked the value layers in deeper locations, even regarded the output generation layers as the value layers, where the knowledge truly be stored in latter ones (details in Appendix F). And the worst is, these KE methods ignored the importance of editing the query layers.
@@
Case Study. As a case study, we examined the residual stream for the single-hop request: “Tim Duncan plays the sport of”. The model exhibited the largest increase in importance at tokens where semantics converge, such as “plays” and “of”, with reaching $0.9932$ and $0.8469$, respectively. The top predicted tokens were highly interpretable (e.g., “basketball”, “ball”, “NBA”). Meanwhile, at other transitional tokens, the model demonstrated stronger exploratory capabilities by predicting incoherent tokens. We select 27 neurons whose associated vocabulary included the correct target token and performed 1000 sampling trials with these neurons removed. The model’s accuracy under this condition dropped to merely 3.2%, indicating the crucial role these interpretable neurons play in generating correct responses. In contrast, when we ablated 27 neurons with high importance scores but lacking correct semantic interpretability, the model maintained 59.4% accuracy. This suggests the final generation of correct output tokens depends on a sparse set of highly specialized, interpretable neurons.We also observe some neurons serve as q-v shared neurons at the same time, which are highly interpretable, the details of this residual stream see in Appendix E.
@@
6.5 Knowledge Bottleneck
In all knowledge editing frameworks based on the locate-then-edit paradigm, the number of critical layers constitutes a crucial parameter. However, extensively editing a large number of model layers proves infeasible and often encounters a knowledge bottleneck, which is particularly pronounced in methods focusing solely on value neurons editing. Therefore, we conduct a sensitivity analysis on the number of edited layers for AcE and other baselines, demonstrating that query neurons editing is both highly efficient and necessary. Since editing query and value neurons incurs equivalent latency in AcE, we begin our comparison of methods’ sensitivity to the knowledge bottleneck with an AcE variant restricted to skipping editing a single query layer, and then we extend the skipping numbers.
As shown in Table 6, we compare the performance of various editors under the condition of an equal number of edited layers, where “#N - editor” denotes the quantity of layers edited by each respective editor. It can be observed that a pronounced knowledge bottleneck exists across all editors except AcE. This manifests specifically as follows: even when significantly increasing the number of edited layers compared to the original baselines, editors focusing solely on value neurons exhibit only marginal performance gains. This phenomenon can be interpreted as evidence of an editing bottleneck inherent to value neurons, wherein excessive modifications fail to enhance the model’s reasoning capability and may even prove detrimental. These results underscore the critical importance of editing query neurons.
Table 6: Analysis of number of edited layers on GPT-J and Qwen3-8B.
Editor Avg. Editor Avg.
#9 - ROME (GPT-J) $37.98$ #8 - ROME (Qwen3-8B) $30.16$
#8 - ROME (GPT-J) $37.56$ #7 - ROME (Qwen3-8B) $29.72$
#7 - ROME (GPT-J) $37.27$ #6 - ROME (Qwen3-8B) $29.52$
#9 - MEMIT (GPT-J) $39.47$ #8 - MEMIT (Qwen3-8B) $21.95$
#8 - MEMIT (GPT-J) $39.14$ #7 - MEMIT (Qwen3-8B) $21.47$
#7 - MEMIT (GPT-J) $38.98$ #6 - MEMIT (Qwen3-8B) $20.16$
#9 - PMET (GPT-J) $38.29$ #8 - PMET (Qwen3-8B) $21.14$
#8 - PMET (GPT-J) $38.16$ #7 - PMET (Qwen3-8B) $21.08$
#7 - PMET (GPT-J) $38.10$ #6 - PMET (Qwen3-8B) $20.92$
#9 - AcE (GPT-J) $46.45$ #8 - AcE (Qwen3-8B) $58.24$
#8 - AcE (GPT-J) $45.28$ #7 - AcE (Qwen3-8B) $57.19$
#7 - AcE (GPT-J) $43.58$ #6 - AcE (Qwen3-8B) $54.10$
@@ 7 Conclusion
Reproducibility Statement
We place strong emphasis on the transparency and reproducibility of our work. We have uploaded the codes we used for reproducibility in the Supplementary Material. To facilitate independent verification, the complete implementation has been provided in the supplementary materials, allowing readers to directly reproduce the reported experiments. In addition, Section 6 and Appendix C, H of the main text outlines the experimental pipeline, including dataset preparation, model configurations, prompts we used and training procedures. For further clarity, Appendix G documents the full set of hyperparameter choices and auxiliary details. Together, these resources ensure that our results can be reliably replicated and extended in future research.
Ethics Statement
This work complies with the ICLR Code of Ethics. All authors of this work have committed to its adherence. The datasets used in this study are publicly available benchmarks. Our research does not involve any private or sensitive personal data. The code developed for experiments will be made publicly available to ensure reproducibility. We have followed standard practices in the field to ensure the fairness and reproducibility of our experiments. Efforts have been made to mitigate potential biases in the evaluation process.
@@
Appendix A Use of Large Language Models
During manuscript preparation, a large language model (LLM) was occasionally employed as an auxiliary assistant to refine language expression, such as improving sentence fluency and enhancing readability. The model was not involved in generating original research contributions: it did not participate in formulating research questions, designing methodologies, conducting experiments, analyzing results, or drafting substantive scientific content. All core intellectual work, including the development of ideas, execution of experiments, and interpretation of findings, was carried out independently by the authors. Any linguistic suggestions offered by the LLM were critically reviewed and selectively incorporated, ensuring that accuracy, originality, and scholarly integrity were fully maintained. The authors alone bear responsibility for the research content and conclusions, and the LLM is not listed as a contributor or author.
Appendix B Related Work in Details
@@
The work by Yu and Ananiadou (2023) provides a crucial foundation for our methodology. They demonstrated that both attention and FFN layers operate via query neurons that activate specific value neurons to produce final predictions. These value neurons were found to be semantically interpretable when projected into the vocabulary space, and similar types of relations (e.g., “capital of X,” “birthplace of Y”) were shown to activate neurons in structurally consistent locations across different subjects. While IFMET focused on attribution and interpretation, AcE leverages these neuron-level insights to develop a more targeted and mechanistically grounded KE strategy, particularly for the complex dynamics of multi-hop reasoning where intermediate implicit subjects act as activators.
@@
Appendix C Subset of MQuAKE Dataset
C.1 Subset Construction
@@
C.2 Dataset Labeling
We use GPT-4o to label the knowledge in the dataset. We show the prompts we used in Appendix H.3. We labeled the dataset related to the knowledge, which are Geographic location, Organization, Personal attributes, Sports, Entertainment, Language and culture, Education, Religion, Literature and Event, which consists of more detailed smaller classes.
Appendix D AcE Optimization
@@
Appendix E Details of Internal Fine-Grained Reasoning Q-V Pairs Flow in Qwen3-8B
@@
Appendix F More Details of the Analysis
@@
Appendix G Experimental Settings
@@
In our edits, our configuration for PMET adheres to the settings specified by (Li et al., 2024). Initially, we set $\varphi=1$ and $0\leq\mu\leq 1$ to manage the retention of the model’s original knowledge. As $\mu$ increases, the retention level also increases, while $\varphi$ exhibits the opposite trend. After maximizing the probability of the target knowledge, we reduce $\varphi$ to 0.1 to preserve the original knowledge as much as possible. Optimization is halted when $D_{KL}<0.01$. On GPT-J and Qwen3-8B, for estimating the covariance matrix (i.e., the set of previously memorized keys $C_{0}$), we sample 10,0000 times on Wikitext in fp32 precision and set $\lambda=6000$. When optimizing, we limit the total optimization steps to 30 with a learning rate of 0.2. All our experiments were conducted using the MQuAKE dataset. To test the accuracy of answers to multi-hop questions, we adhered to the few-shot and Chain of Thought (CoT) templates in Appendix H.1 and procedures.
Appendix H Prompts
@@
H.1 Few-shot and Chain-of-Thought Evaluation Prompts